# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 5.5 Cyclotron Damping
For the fundamental ($n=1$) cyclotron resonance, the resonance condition in (20) simplifies to $(\omega-k_{\parallel}v_{\parallel})/\omega_{s}=1$. The numerator involves the fluctuation frequency modified by the Doppler shift of the particles motion through spatial variations parallel to the magnetic field. When that Doppler-shifted frequency matches the ion cyclotron frequency, the perpendicular compnent of the electric field $E_{\perp}$ can resonantly interact with the particles to energize the particles in the directions perpendicular to the magnetic field, yielding $Q_{\perp,s}/Q_{s}=1$. To estimate the conditions required for cyclotron resonant damping, it is often sufficient to require the condition $\omega/\omega_{s}\rightarrow 1$.
In typical space and astrophysical plasmas, under the conditions derived below the turbulent cascade can reach the ion cyclotron frequency $\omega/\omega_{i}\rightarrow 1$, but since the electron cyclotron frequency is higher by the mass ratio, $\Omega_{e}/\Omega_{i}=m_{i}/m_{e}$, there is typically little energy remaining in the cascade at the very high frequencies needed to transfer is significant amount of energy to the electron through their cyclotron resonance. Thus, here we consider only the physics of the ion cyclotron resonance as a possible physical mechanism for the dissioation of plasma turbulence.
@@
where $\overline{\omega}$ is defined by (25). Note that the middle expression indicates that the most natural unit to normalize the length scale parallel to the magnetic field is the ion inertial length $d_{i}=c/\omega_{pi}=v_{A}/\Omega_{i}$, but to maintain consistency with the length normalizations in Tables 2 and  3, we substitute $d_{i}=\rho_{i}\beta_{i}^{-1/2}$. Utilizing the general scaling of the parallel wavenumber in terms of the isotropic driving wavenumber,
@@
This indicates clearly that ion cyclotron damping is more likely to occur for small ion plasma beta $\beta_{i}\ll 1$ or a smaller turbulent inertial range with a value of $k_{0}\rho_{i}$ this is not too small. For the GS95 scaling with $\alpha=0$, we obtain a condition for the onset of ion cyclotron damping of $k_{0}\rho_{i}=\beta_{i}^{3/2}/\overline{\omega}^{3}$; for the B06 scaling with $\alpha=1$, we obtain $k_{0}\rho_{i}=\beta_{i}^{/2}/\overline{\omega}^{2}$. For simplicity, note that at the small-scale end of the inertial range with $k_{\perp}\rho_{i}\sim 1$, the normalized wave frequency $\overline{\omega}\sim 1$.
In summary, ion cyclotron damping is predicted to play a role in the damping of plasma turbulence for conditions with a smaller inertial range and a low plasma beta values $\beta_{i}\ll 1$. The nature of the resulting plasma heating is strictly ion energization $Q_{i}/Q=1$ in the perpendicular degrees of freedom $Q_{\perp,i}/Q_{i}=1$.
Note that recent theoretical and numerical study of turbulence with a sufficiently large amount of imbalance $Z_{0}^{+}/Z_{0}^{-}\gg 1$, such as may be encountered in the inner heliosphere, has uncovered a new phenomenon denoted the helicity barrier, which can lead to the generation of sufficiently high-frequency fluctuations, specifically ion cyclotron waves, to lead to iCD energization of the plasma particles (Squire et al., 2022, 2023).
@@ 5.6 Stochastic Heating
The stochastic heating of ions has long been proposed as a means to energize ions in turbulent plasmas (Johnson & Cheng, 2001; Chen et al., 2001; White et al., 2002; Voitenko & Goossens, 2004; Bourouaine et al., 2008), with more recent work providing significant theoretical and numerical evidence for ion stochastic heating as a significant mechanism for turbulent damping in space and astrophysical plasmas (Chandran et al., 2010; Chandran, 2010; Chandran et al., 2011; Bourouaine & Chandran, 2013; Chandran et al., 2013; Klein & Chandran, 2016; Arzamasskiy et al., 2019; Martinović et al., 2020; Cerri et al., 2021).
Chandran et al. (2010) used test particle simulations of randomly phased Alfvén and kinetic Alfvén waves to develop an empirical formula for the rate of ion stochastic heating by turbulent fluctuations with frequencies much lower than the ion cyclotron frequency $\omega/\Omega_{i}\ll 1$. Following this, Chandran et al. (2011) developed a refined implementation to predict the partitioning of turbulent energy among parallel ion heating, perpendicular ion heating, and electron heating, with a formula for the damping rate by ion stochastic heating given by
@@
where $\delta v_{\perp}^{(\rho_{i})}$ is the root-mean-square amplitude of the perpendicular velocity fluctuations at the perpendicular scale of the ion Larmor radius $k_{\perp}\rho_{i}=1$. Here $c_{1}$ and $c_{2}$ are fitting constants with typical values $c_{1}=0.18$ and $0.15\lesssim c_{2}\lesssim 0.34$ (Chandran et al., 2011). The key dimensionless parameter controlling the onset of stochastic heating in this formulation is the ratio of the perpendicular velocity fluctuation to the thermal velocity, $\epsilon\equiv\delta v_{\perp}^{(\rho_{i})}/v_{ti}$; these studies found a critical threshold $\epsilon_{crit}\simeq 0.19$ that determines the minimum fluctuation amplitude required for the onset of stochastic heating (Chandran et al., 2010).
@@
where we have used the subsitution $\Omega_{i}/\omega=\beta_{i}^{1/2}/k_{\parallel}\rho_{i}$.
@@
Note that, because the amplitude of the turbulent fluctuations is typically found to decrease monotonically as the turbulent cascade progress to smaller scales, it is unlikely that electrons will experience significant stochastic heating, so we do not explore electron stochastic heating here. Further research will be necessry to to evaluate thoroughly the potential for electron stochastic heating to damp turbulence at the perpendicular scale of the electron Larmor radius, $k_{\perp}\rho_{e}\sim 1$.
@@ 5.7 Magnetic Pumping
Magnetic pumping (Spitzer & Witten, 1951; Berger et al., 1958; Lichko et al., 2017; Lichko & Egedal, 2020; Montag & Howes, 2022) is a particle energization mechanism in which variations of magnetic field magnitude at low frequency $\omega$ lead to an oscillating transfer of energy from parallel to perpendicular degrees of freedom and back, obeying the double-adiabatic (Chew-Goldberger-Low, CGL) equations of state (Chew et al., 1956). In the absence of any collisions, the net change of particle energy over a full oscillation is zero. But in the presence small but finite collisionality [^9] Note that the collisionality can be due to Coulomb collisions or to an effective collisionality due to pitch-angle scattering through collisionless wave-particle interactions with small-scale electromagnetic fluctuations, such as those driven by kinetic temperature anisotropy instabilities (Kunz et al., 2018; Arzamasskiy et al., 2023). there can be a non-zero transfer of energy from the oscillating electromagnetic fields to the particles over the course of a full oscillation.
As a proposed mechanism for the damping of turbulence, magnetic pumping necessarily requires turbulent fluctuations that involve changes of the magnetic field magnitude from the equilibrium field strength $B_{0}$, or $\delta B\equiv|\mathbf{B}|-B_{0}\neq 0$. For Alfvénic turbulence, which is dominantly incompressible [^10] Note, however, that the magnetic field magnitude variations arising from strong turbulent driving at the outer scale with $|\delta\mathbf{B}|\sim B_{0}$ can lead to the “interruption” of linearly polarized Alfvén waves by kinetic temperature anisotropy instabilities at sufficient high $\beta_{i}\gg 1$ (Squire et al., 2016, 2017b, 2017a) in the MHD regime at $k_{\perp}\rho_{i}\ll 1$, significant fluctuations of the magnetic field magnitude arise only within the kinetic Alfvén wave regime at $k_{\perp}\rho_{i}\gtrsim 1$ under conditions of ion plasma beta $\beta_{i}>1$. Alternatively, if the turbulence has a signficant component of compressible magnetosonic fluctuations with $E_{comp}/E_{inc}>0$, because the turbulent amplitudes typically decrease monotonically with decreasing scale lengths, the largest scale compressible fluctuations at the driving scale may lead to magnetic pumping, directly energizing particles from these large-scale fluctuations.
@@
Kinetic Alfvén waves give rise to non-negligible magnetic field magnitude variations over the perpendicular wavenumber range $1\lesssim k_{\perp}\rho_{i}\lesssim(\tau\mu)^{1/2}$ for values of $\beta_{i}\gtrsim 1$. This range of $\delta B$ fluctuations is incompatible with the necessary condition $k_{\perp}\rho_{i}\ll 1$ for the ions, so we expect that magnetic pumping by kinetic Alfvén waves cannot energize the ions. For the electrons, on the other hand, the requisite condition $k_{\perp}\rho_{e}\ll 1$ can be rewritten as $k_{\perp}\rho_{i}\ll(\tau\mu)^{1/2}$, so we expect that kinetic Alfvén waves may be able to damp turbulent fluctuations and energize the electrons under ion plasma beta conditions with $\beta_{i}\gtrsim 1$. The phase velocity condition for electrons can be expressed as $\overline{\omega}\ll(\beta_{i}\mu/\tau)^{1/2}$. Under the ion plasma beta condition $\beta_{i}\gtrsim 1$, (25) can be approximated by $\overline{\omega}\simeq k_{\perp}\rho_{i}/\beta_{i}^{1/2}$, so this requirement can be converted to a condition on the perpendicular wavenumber given by $k_{\perp}\rho_{i}\ll\beta_{i}(\mu/\tau)^{1/2}$, which can certainly be satisfied for the perpendicular wavenumber range of kinetic Alfvén waves with significant magnetic field magnitude fluctuations. Therefore, the necessary conditions for the magnetic pumping of electrons to play a role in the dissipation of Alfvénic plasma turbulence with $\beta_{i}\gtrsim 1$ appear to be satisfied.
@@
To determine the dependence of the magnetic pumping of electrons on the plasma and turbulence parameters in Tables 2 and  3, we assume isotropic driving $k_{\parallel 0}/k_{\perp 0}=1$ for simplicity, write the electron-electron collision frequency as $\nu_{e}=v_{te}/\lambda_{mfp,e}$, and use the general parallel wavenumber scaling for arbitrary $\alpha$ given by (1) to obtain
@@
$\widetilde{\omega}^{4}-\widetilde{\omega}^{2}(1+\beta)+\beta\cos^{2}\Theta_{0}=0,$ (36)
@@
$\widetilde{\omega}^{2}=\frac{1}{2}\left[1+\beta\pm\sqrt{(1+\beta)^{2}-4\beta\cos^{2}\Theta_{0}}\right],$ (37)
@@
To evaluate the turbulent magnetic field magnitude variations for ion magnetic pumping by compressible fluctuations at the driving scale, we first note that the phenomenon of dynamic alignment, parameterizied by the alignment angle $\theta_{0}$, will not play a role: dynamic alignment develops through the nonlinear interactions as the turbulence cascades to smaller scales, but does not come into play at the driving scale, so we set $\theta_{0}=1$, thus reducing the nonlinearity parameter to $\chi_{0}=k_{\perp 0}\delta B_{\perp 0}/(k_{\parallel 0}B_{0})$. If the magnetic perturbations at the driving scale have $|\delta\mathbf{B}|/B_{0}\sim 1$, the linear theory for waves breaks down, and the nonlinearity must be taken into account in the lowest-order description of the wave dynamics; in this case, one may directly take the measured magnetic field magnitude fluctuations $\delta B/B_{0}$ as the parameter to estimate the damping rate by magnetic pumping in (33). However, if the magnetic perturbations at the driving scale have $\epsilon\equiv|\delta\mathbf{B}|/B_{0}<1$, the binomial expansion may be used to obtain the expression
@@
In summary, Alfvénic turbulent fluctuations can damping turbulence and energize electrons through turbulent fluctuations at perpendicular scales $1\lesssim k_{\perp}\rho_{i}\lesssim(\tau\mu)^{1/2}$ for plasma conditions with $\beta_{i}\gtrsim 1$, but ions are not expected to be energized by Alfvénic fluctuations. The resulting magnetic pumping of electrons by Alfvénic turbulent fluctuations depends on the parameters $\beta_{i}$, $\tau$, $\mu$, $k_{0}\rho_{i}$, and $k_{\parallel 0}\lambda_{mfp,e}$. If the turbulence is driven compressibly with $E_{comp}/E_{inc}>0$, we expect that the turbulent fluctuations at the driving scales, which have the largest amplitude of $\delta B/B_{0}$, may lead to damping of the turbulence through the magnetic pumping of both the ions and electrons. The resulting magnetic pumping of ions by compressible turbulence at the driving scales depends on the parameters $\beta_{i}$, $k_{\parallel 0}\lambda_{mfp,i}=(k_{\parallel 0}\lambda_{mfp,e})\tau^{2}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$, and $E_{comp}/E_{inc}$. Analogously, the magnetic pumping of electrons by compressible turbulence depends on the parameters $\beta_{i}$, $\tau$, $\mu$, $k_{\parallel 0}\lambda_{mfp,e}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$, and $E_{comp}/E_{inc}$. These parameter dependencies for magnetic pumping of ions and electrons are summarized in Table 4.

### v2

@@ 5.5 Cyclotron Damping
For the fundamental ($n=1$) cyclotron resonance, the resonance condition in (20) simplifies to $(\omega-k_{\parallel}v_{\parallel})/\Omega_{s}=1$. The numerator involves the fluctuation frequency modified by the Doppler shift of the particle’s motion parallel to the magnetic field. When that Doppler-shifted frequency matches the ion cyclotron frequency, the perpendicular component of the electric field $E_{\perp}$ can resonantly interact with the particles to energize the particles in the directions perpendicular to the magnetic field, yielding $Q_{\perp,s}/Q_{s}=1$. To estimate the conditions required for cyclotron resonant damping, it is often sufficient to require the condition $\omega/\Omega_{s}\rightarrow 1$.
In typical space and astrophysical plasmas, the turbulent cascade can reach the ion cyclotron frequency $\omega/\Omega_{i}\rightarrow 1$ under the conditions derived below, but since the electron cyclotron frequency is higher by the mass ratio, $\Omega_{e}/\Omega_{i}=m_{i}/m_{e}$, there is typically little energy remaining in the cascade at the very high frequencies needed to transfer a significant amount of energy to the electrons through their cyclotron resonance. Thus, here we consider only the physics of the ion cyclotron resonance as a possible physical mechanism for the dissipation of plasma turbulence.
@@
where $\overline{\omega}$ is defined by (25). Note that the middle expression indicates that the most natural unit to normalize the length scale parallel to the magnetic field is the ion inertial length $d_{i}=c/\omega_{pi}=v_{A}/\Omega_{i}$, but to maintain consistency with the length normalizations in Tables 3 and  4, we substitute $d_{i}=\rho_{i}/\beta_{i}^{1/2}$. Utilizing the general scaling of the parallel wavenumber in terms of the isotropic driving wavenumber,
@@
This indicates clearly that ion cyclotron damping is more likely to occur for small ion plasma beta $\beta_{i}\ll 1$ or a smaller turbulent inertial range (meaning a value of $k_{0}\rho_{i}$ that is not too small). For the GS95 scaling with $\alpha=0$, we obtain a condition for the onset of ion cyclotron damping of $k_{0}\rho_{i}=\beta_{i}^{3/2}/\overline{\omega}^{3}$; for the B06 scaling with $\alpha=1$, we obtain $k_{0}\rho_{i}=\beta_{i}/\overline{\omega}^{2}$. For simplicity, note that at the small-scale end of the inertial range with $k_{\perp}\rho_{i}\lesssim 1$, the normalized wave frequency $\overline{\omega}\sim 1$ for Alfvén waves can be approximated in the limit $k_{\parallel}d_{i}\rightarrow 1$ by the expression (Di Mare & Howes, 2024)
$\overline{\omega}\simeq\frac{1}{\sqrt{1+(k_{\parallel}d_{i})^{2}}}.$ (30)
In summary, ion cyclotron damping is predicted to play a role in the damping of plasma turbulence for conditions with a smaller inertial range and low plasma beta values $\beta_{i}\ll 1$. The nature of the resulting plasma heating is strictly ion energization $Q_{i}/Q=1$ in the perpendicular degrees of freedom $Q_{\perp,i}/Q_{i}=1$.
Note that recent theoretical and numerical studies of turbulence with a sufficiently large amount of imbalance $Z_{0}^{+}/Z_{0}^{-}\gg 1$, such as may be encountered in the inner heliosphere, has uncovered a new phenomenon denoted the helicity barrier, which can lead to the generation of sufficiently high-frequency fluctuations, specifically ion cyclotron waves, that may cause energization of the ions by the cyclotron resonance (Meyrand et al., 2021; Squire et al., 2022, 2023).
@@ 5.6 Stochastic Heating
The stochastic heating of ions has long been proposed as a means to energize ions in turbulent plasmas (Johnson & Cheng, 2001; Chen et al., 2001; White et al., 2002; Voitenko & Goossens, 2004; Bourouaine et al., 2008), with more recent work providing significant theoretical and numerical evidence for ion stochastic heating as a significant mechanism for turbulent damping in space and astrophysical plasmas (Chandran et al., 2010; Chandran, 2010; Chandran et al., 2011; Bourouaine & Chandran, 2013; Chandran et al., 2013; Klein & Chandran, 2016; Vech et al., 2017; Arzamasskiy et al., 2019; Martinović et al., 2020; Cerri et al., 2021).
Chandran et al. (2010) used test particle simulations of randomly phased Alfvén and kinetic Alfvén waves to develop an empirical formula for the rate of ion stochastic heating by turbulent fluctuations with frequencies much lower than the ion cyclotron frequency $\omega/\Omega_{i}\ll 1$. Following this, Chandran et al. (2011) developed a refined implementation to predict the partitioning of turbulent energy among parallel ion heating, perpendicular ion heating, and electron heating (the C11 model in §4), with a formula for the damping rate by ion stochastic heating given by
@@
where $\delta v_{\perp}^{(\rho_{i})}$ is the root-mean-square amplitude of the perpendicular velocity fluctuations at the perpendicular scale of the ion Larmor radius, $k_{\perp}\rho_{i}=1$. Here $c_{1}$ and $c_{2}$ are fitting constants with suggested values $c_{1}=0.18$ and $0.15\lesssim c_{2}\lesssim 0.34$ (Chandran et al., 2011). The key dimensionless parameter controlling the onset of stochastic heating in this formulation is the ratio of the perpendicular velocity fluctuation to the thermal velocity, $\epsilon\equiv\delta v_{\perp}^{(\rho_{i})}/v_{ti}$; these studies found a critical threshold $\epsilon_{crit}\simeq 0.19$ that determines the minimum fluctuation amplitude required for the onset of stochastic heating (Chandran et al., 2010).
@@
where we have used the substitution $\Omega_{i}/\omega=\beta_{i}^{1/2}/k_{\parallel}\rho_{i}$.
@@
Note that, because the amplitude of the turbulent fluctuations is typically found to decrease monotonically as the turbulent cascade progress to smaller scales, it is unlikely that electrons will experience significant stochastic heating, so we do not explore electron stochastic heating here. Further research will be necessary to evaluate thoroughly the potential for electron stochastic heating to damp turbulence at the perpendicular scale of the electron Larmor radius, $k_{\perp}\rho_{e}\sim 1$. Furthermore, a consideration of the intermittency—which can enhance the amplitude of individual fluctuations—of the turbulent fluctuations at the ion Larmor radius scale suggests that the rate of ion stochastic heating may be dramatically increased in the presence of significant intermittency (Mallet et al., 2019).
@@ 5.7 Magnetic Pumping
Magnetic pumping (Spitzer & Witten, 1951; Berger et al., 1958; Lichko et al., 2017; Lichko & Egedal, 2020; Montag & Howes, 2022) is a particle energization mechanism in which variations of magnetic field magnitude at low frequency $\omega$ lead to an oscillating transfer of energy from parallel to perpendicular degrees of freedom and back, obeying the double-adiabatic (Chew-Goldberger-Low, CGL) equations of state (Chew et al., 1956). In the absence of any collisions, the net change of particle energy over a full oscillation is zero. But, in the presence small but finite collisionality [^12] Note that the collisionality can be due to Coulomb collisions or to an effective collisionality due to pitch-angle scattering through collisionless wave-particle interactions with small-scale electromagnetic fluctuations, such as those driven by kinetic temperature anisotropy instabilities (Kunz et al., 2018; Arzamasskiy et al., 2023)., there can be a non-zero transfer of energy from the oscillating electromagnetic fields to the particles over the course of a full oscillation.
As a proposed mechanism for the damping of turbulence, magnetic pumping necessarily requires turbulent fluctuations that involve changes of the magnetic field magnitude from the equilibrium field strength $B_{0}$, or $\delta B\equiv|\mathbf{B}|-B_{0}\neq 0$. For Alfvénic turbulence, which is dominantly incompressible [^13] Note, however, that the magnetic field magnitude variations arising from strong turbulent driving at the outer scale with $|\delta\mathbf{B}|\sim B_{0}$ can lead to the “interruption” of linearly polarized Alfvén waves by kinetic temperature anisotropy instabilities at sufficient high $\beta_{i}\gg 1$ (Squire et al., 2016, 2017a, 2017b) in the MHD regime at $k_{\perp}\rho_{i}\ll 1$, significant fluctuations of the magnetic field magnitude arise only within the kinetic Alfvén wave regime at $k_{\perp}\rho_{i}\gtrsim 1$ under conditions of ion plasma beta $\beta_{i}>1$. Alternatively, if the turbulence has a significant component of compressible magnetosonic fluctuations with $E_{comp}/E_{inc}>0$, because the turbulent amplitudes typically decrease monotonically with decreasing scale lengths, the largest amplitude compressible fluctuations occur at the driving scale and may lead to magnetic pumping, directly energizing particles from these large-scale fluctuations.
@@
Kinetic Alfvén waves give rise to non-negligible magnetic field magnitude variations over the perpendicular wavenumber range $1\lesssim k_{\perp}\rho_{i}\lesssim(\tau\mu)^{1/2}$ for values of $\beta_{i}\gtrsim 1$. This range of $\delta B$ fluctuations is incompatible with the necessary condition $k_{\perp}\rho_{i}\ll 1$ for the ions, so we expect that magnetic pumping by kinetic Alfvén waves cannot energize the ions. For the electrons, on the other hand, the requisite condition $k_{\perp}\rho_{e}\ll 1$ can be rewritten as $k_{\perp}\rho_{i}\ll(\tau\mu)^{1/2}$, so we expect that kinetic Alfvén waves may be able to damp turbulent fluctuations and energize the electrons under ion plasma beta conditions with $\beta_{i}\gtrsim 1$. The phase velocity condition for electrons can be expressed as $\overline{\omega}\ll(\beta_{i}\mu/\tau)^{1/2}$. Under the ion plasma beta condition $\beta_{i}\gtrsim 1$, the equation for the dimensionless frequency of kinetic Alfvén waves given by (26) can be approximated by $\overline{\omega}\simeq k_{\perp}\rho_{i}/\beta_{i}^{1/2}$, so this requirement can be converted to a condition on the perpendicular wavenumber given by $k_{\perp}\rho_{i}\ll\beta_{i}(\mu/\tau)^{1/2}$, which can certainly be satisfied for the perpendicular wavenumber range of kinetic Alfvén waves with significant magnetic field magnitude fluctuations. Therefore, the necessary conditions for the magnetic pumping of electrons to play a role in the dissipation of Alfvénic plasma turbulence with $\beta_{i}\gtrsim 1$ appear to be satisfied.
@@
To determine the dependence of the magnetic pumping of electrons on the plasma and turbulence parameters in Tables 3 and  4, we assume isotropic driving $k_{\parallel 0}/k_{\perp 0}=1$ for simplicity, write the electron-electron collision frequency as $\nu_{e}=v_{te}/\lambda_{mfp,e}$, and use the general parallel wavenumber scaling for arbitrary $\alpha$ given by (1) to obtain
@@
$\widetilde{\omega}^{4}-\widetilde{\omega}^{2}(1+\beta)+\beta\cos^{2}\Theta=0,$ (38)
@@
$\widetilde{\omega}^{2}=\frac{1}{2}\left[1+\beta\pm\sqrt{(1+\beta)^{2}-4\beta\cos^{2}\Theta}\right],$ (39)
@@
To evaluate the turbulent magnetic field magnitude variations for ion magnetic pumping by compressible fluctuations at the driving scale, we first note that the phenomenon of dynamic alignment, parameterized by the alignment angle $\theta_{0}$, will not play a role: dynamic alignment develops through the nonlinear interactions as the turbulence cascades to smaller scales, but does not come into play at the driving scale, so we set $\theta_{0}=1$, thus reducing the nonlinearity parameter at the driving scale to $\chi_{0}=k_{\perp 0}\delta B_{\perp 0}/(k_{\parallel 0}B_{0})$. If the magnetic perturbations at the driving scale have $|\delta\mathbf{B}|/B_{0}\sim 1$, the linear theory for waves breaks down, and the nonlinearity must be taken into account in the lowest-order description of the wave dynamics; in this case, one may directly take the measured magnetic field magnitude fluctuations $\delta B/B_{0}$ as the parameter to estimate the damping rate by magnetic pumping in (35). However, if the magnetic perturbations at the driving scale have $\epsilon\equiv|\delta\mathbf{B}|/B_{0}<1$, the binomial expansion may be used to obtain the expression
@@
In summary, Alfvénic turbulent fluctuations can damp turbulence and energize electrons through turbulent fluctuations at perpendicular scales $1\lesssim k_{\perp}\rho_{i}\lesssim(\tau\mu)^{1/2}$ for plasma conditions with $\beta_{i}\gtrsim 1$, but ions are not expected to be energized by Alfvénic fluctuations. The resulting magnetic pumping of electrons by Alfvénic turbulent fluctuations depends on the parameters $\beta_{i}$, $\tau$, $\mu$, $k_{0}\rho_{i}$, and $k_{\parallel 0}\lambda_{mfp,e}$. If the turbulence is driven compressibly with $E_{comp}/E_{inc}>0$, we expect that the turbulent fluctuations at the driving scales, which have the largest amplitude of $\delta B/B_{0}$, may lead to damping of the turbulence through the magnetic pumping of both the ions and electrons. The resulting magnetic pumping of ions by compressible turbulence at the driving scales depends on the parameters $\beta_{i}$, $k_{\parallel 0}\lambda_{mfp,i}=(k_{\parallel 0}\lambda_{mfp,e})\tau^{2}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$, and $E_{comp}/E_{inc}$. Analogously, the magnetic pumping of electrons by compressible turbulence depends on the parameters $\beta_{i}$, $\tau$, $\mu$, $k_{\parallel 0}\lambda_{mfp,e}$, $k_{\parallel 0}/k_{\perp 0}$, $\chi_{0}$, and $E_{comp}/E_{inc}$. These parameter dependencies for magnetic pumping of ions and electrons are summarized in Table 5. Note that magnetic pumping energizes particles in energy over all pitch angles (Montag & Howes, 2022), so that both the parallel and perpendicular degrees of freedom gain energy, yielding for either ion or electron species $Q_{\perp,s}/Q_{\parallel,s}\sim 1$.